Suppose that you and a friend are playing cards and you decide to make a friendly wager. The bet is that you will draw two cards without replacement from a standard deck. If both cards are clubs, your friend will pay you $9. Otherwise, you have to pay your friend $3. Step 2 of 2: If this same bet is made 753 times, how much would you expect to win or lose? Round your answer to two decimal places. Losses must be expressed as negative values.
Answer #1

Ans:

Probability of drawing two club cards without replacement

=(13/52)*(12/51)=0.0588

Expected loss on single bet=0.0588*9-3*(1-0.0588)=-2.29

If repeated 753 times:

Expected loss=-2.29*753=-1727.47